# Defensive Prompt Patch: A Robust and Interpretable Defense of LLMs against Jailbreak Attacks

### v1

@@ 3.3 DPP Training Algorithm
10:         Find synonyms for each word
11:         if random value $<$ WordDict[$synonym$] / sum($word~{}scores$then
12:              Replace word with synonym
13:         end if
14:     end for
15:     for $iteration$ in paragraph-level iterations do
16:         Repeat line 6 to 8
17:         Conduct crossover and mutation on selected parent prompts using Alg. 4
18:     end for
@@
21:end while
22:return (New_DPP_Set, Best_DPP)
@@
Algorithm 1 identifies the optimal DPP for a given pair of refusal and helpful data. Our primary objective is to find a DPP that generalizes well across different user queries. To enhance the universality of DPP, we incorporate $N$ pairs of refusal and helpful data, sampled from their respective datasets. In each iteration of the DPP algorithm, as described earlier, a set of candidate DPPs is generated along with the best DPP for the specific data pair. This set of candidate DPPs is then used for the next pair of refusal and helpful data. By iteratively optimizing this set of DPP candidates, we aim to identify the most generalizable DPP with the best defensive and utility performance. The overall optimization procedure is detailed in Algorithm 5. For full implementation details and hyperparameter settings, please refer to Appendix D.
@@ 4 Experiments
We demonstrate the performance of our DPP through three perspectives: Robustness to standard (non-adaptive) and adaptive jailbreak attacks, Generalization to unforeseen jailbreak queries and different LLMs, and Interpretability of the best-found DPP. All final DPPs are listed in Appendix H.
@@ 4.1 Experimental Setup
Adversarial Dataset: We use the AdvBench [1], specifically the harmful behavior instructions  [^1] https://github.com/llm-attacks/llm-attacks/blob/main/data/advbench/harmful_behaviors.csv, as jailbreak questions. Each of them is fed into a well-aligned LM (LLAMA-2-7B-Chat [2]) to generate the denial responses. In our experiment, we sampled 100 jailbreak questions and recorded both jailbreak questions along with their refusal responses to form the Adversarial Dataset.
@@
Language Models: We perform our jailbreak experiments on two specific LLMs: LLAMA-2-7B-Chat [2] and Mistral-7B-Instruct-v0.2 [7]. LLAMA-2-7B-Chat is an adapted version of LLAMA-2-7B, specifically configured for chat-based interactions. Mistral-7B-Instruct-v0.2 is a fine-tuned chat version of Mistral-7B-v0.2. This model demonstrates a stronger ability in performance, outperforming LLAMA-2-13B on all benchmarks while maintaining proficiency in English language tasks.
@@
Jailbreak Attack Methods: We use several existing jailbreak attack methods to generate advanced malicious prompts. Specifically, for each malicious behavior statement, we apply several different types of jailbreaking attacks: (i) Uninterpretable Jailbreak Attacks – we used GCG [1] and Base64 [6] to generate adversarial prompts. Specifically, GCG is used to generate an adversarial suffix for each malicious query. Base64 encodes each harmful query in Base64 format. (ii) Interpretable Jailbreak Attacks – AutoDAN [3], PAIR [5], TAP [4], and ICA [8] are interpretable attacks that we used to translate the original malicious query into a new improved malicious query. Please refer to Appendix A for more details on generating new malicious queries. (iii) Generation-based Jailbreak Attacks – we follow Catastrophic Attack [9] to vary the hyperparameters of the LLM to generate malicious responses for each harmful question. In our evaluation, similar to the Adversarial Dataset, we utilize 100 harmful behavior questions from AdvBench to generate new malicious queries [^3] For PAIR and TAP adaptive attacks, we directly utilize the dataset provided in their code-base, which they sample 50 harmful behaviors from AdvBench., all of which will be employed in our experiments.
@@
Jailbreak Defense Methods: We compare our DPP to Self-Reminder [10] and Goal Prioritization [11]. They are prompt-based defenses that add defense prompts as a prefix or suffix. For the LLAMA-2-7B chat model, we also include another defensive suffix approach called RPO [12]. For Mistral-7B-Instruct-v0.2, instead of using RPO as a baseline, we compare the results with Plain (Default) System Prompt [15]. We defer the discussion of our choices of baselines for the two LLMs to Appendix B. The prompts for each defense can be found in Appendix G.
@@
Evaluation Metrics: We use the Attack Success Rate (ASR) as our primary metric for evaluating the effectiveness of jailbreak defenses. The ASR measures the proportion of malicious queries that successfully bypass the LLMs alignment and generate harmful responses. Details on how we calculate ASR can be found in Appendix  C. In addition to ASR, we also use AlpacaEval [14] to evaluate the utility degradation of the LLM model when defenses are employed. Specifically, we utilize the metric called Win-Rate. This involves comparing the frequency with which outputs from LLM are favored over those from a reference model, given a specific user instruction. Utilizing simulated Win-Rate offers a straightforward, comparable metric across various LLMs using the same reference model. In Appendix O, we discuss the setups of evaluating with Win-Rate.
@@ 4.2 Robustness against Non-adaptive and Adaptive Attacks
RPO [12] 0.000 0.420 0.280 0.190 0.060 0.060 0.168 79.23
Goal Priorization [11] 0.000 0.020 0.520 0.020 0.020 0.020 0.100 34.29
Self-Reminder [10] 0.030 0.290 0.000 0.040 0.020 0.000 0.063 64.84
@@
Adaptive Methods ICA [$\downarrow$] Catastrophic [$\downarrow$] GCG [$\downarrow$] AutoDAN [$\downarrow$] Average Adaptive ASR [$\downarrow$]
Self-Reminder 0.410 0.263 0.210 0.080 0.241
RPO 0.360 0.653 0.920 0.170 0.526
Goal Prioritization 0.660 0.0033 0.190 0.530 0.346
DPP (Ours) 0.160 0.247 0.120 0.110 0.159
Table 3: Adaptive Attack Success Rates Rate on LLAMA-2-7B-Chat model. Our method can achieve the lowest Average Adaptive ASR.
@@
Adaptive Attacks: Adaptive attack [16] is a critical evaluation procedure for assessing defense effectiveness when the defense mechanism is known to the attack. Here, we assume the attacker can query the protected LLM with the defense in place when making jailbreak attempts. In this setup, we adapted the attack strategies described in Appendix I. Due to the known limited effectiveness of PAIR and TAP in the non-adaptive setting on the LLAMA-2-7B-Chat model, [5, 4], we replace these attacks with Catastrophic Adaptive Attack. In addition, Base64 attack is a static approach, so the adaptive setting cannot be directly applied to it. Therefore, we remove these attacks from the evaluation. Table 3 shows the adaptive attack results. Our method still has the best adaptive ASR with respect to ICA and GCG adaptive attacks. Although Goal Prioritization has the best ASR under catastrophic attacks, which is 0.33%, it fails to defend against ICA and AutoDAN adaptive attacks. On the other hand, our method outperforms Self-Reminder against all adaptive attacks except AutoDAN. Notably, our method attains the best Average ASR, which is 15.9% (outperforming the second-best method by more than 8%), while RPO has the worst robustness, with an Average ASR of 52.6%. In Appendix F, we also conducted our DPP with different initialized prototypes and found the defensive performance was consistent.
In conclusion, both non-adaptive and adaptive evaluations affirm that our DPP consistently surpasses other defense mechanisms in robustness, with minimal utility degradation across the board. This comprehensive performance solidifies our method’s position as a preferable choice for defending the LLAMA-2-7B-Chat model against diverse and sophisticated attacks.
@@ 4.3 Generalization of DPP
Self-Reminder [10] 0.550 0.270 0.510 0.880 0.420 0.260 0.482 88.82
System Prompt [15] 0.740 0.470 0.300 0.970 0.500 0.180 0.527 84.97
Goal Priorization [11] 0.030 0.440 0.030 0.390 0.300 0.140 0.222 56.59
@@
Table 4: Attack Success Rates (ASRs) and Win-Rates (utility) on Mistral-7B-Instruct-v0.2 model across six different jailbreak attacks. Our method can achieve the lowest Average attack success rate with reasonable trade-off of Win-Rate when compared with other defense baselines.
@@
Table 5: Adaptive Attack Success Rates on Mistral-7B-Instruct-v0.2. Our method can achieve the lowest Average ASR.
@@
Regarding the trade-off between defense effectiveness and utility degradation, unlike the LLAMA-2-7B-Chat results, our method exhibits a higher utility degradation, as indicated by the Win-Rate, compared to Self-Reminder, and System Prompt. Nonetheless, the superior defense performance (a gap greater than 46% in average ASR) of our method justifies this increased utility degradation. It is noteworthy that despite the relatively higher utility impact, our method still shows much less degradation compared to the Goal Prioritization approach. Our result suggests that Mistral-7B-Instruct-v0.2 has a worse defense-utility trade-off than LLAMA-2-7B-Chat. That is, the cost of making Mistral-7B-Instruct-v0.2 robust to jailbreak attacks on utility is more significant than LLAMA-2-7B-Chat. We present additional experiments in Appendix P, where we compare our results with another defense baseline and observe similar effects.
Adaptive Attacks: Table 5 demonstrates that our method consistently performs best as a defense mechanism against jailbreak attacks on average. Although our approach is slightly less effective in the GCG Adaptive Attack compared to Goal Prioritization, it exhibits superior defensive capabilities in the AutoDAN, PAIR, and TAP adaptive attacks.
Unforeseen Jailbreak Queries: We also test the generalization of each defense using the JailbreakBench Chat dataset (JBC) [33], which contains harmful queries distinct from those found in the AdvBench dataset. The results from Table 12 in Appendix L show that for the well-aligned model (LLAMA-2-7B-Chat), the JBC dataset does not yield effective jailbreak attacks, resulting in comparable defense performances across all methods. Conversely, with the less-aligned Mistral-7B-Instruct-v0.2 model, our DPP demonstrated its efficacy by reducing the Attack Success Rate (ASR) from 41% to 1%, attaining the best defense performance (on par with Goal Prioritization). This marked decrease in ASR highlights our DPP’s strong capability to generalize defense performance effectively against unforeseen attacks.
@@
4.4 Interpretability of DPP
We explore the interpretability of our DPP by presenting our best DPPs trained on both LLAMA-2-7B-Chat and Mistral-7B-Instruct-v0.2 models below.
DPP (Suffix) for LLAMA-2-7B-Chat Kindly furnish a thorough response to the former user’s question.
We observe that both DPPs are fluent and relevant to the purpose of alignment. One interesting finding is that the optimized DPP for Mistral-7B-Instruct-v0.2 is very explicit in instructing refusal when facing “defective components”, whereas the DPP of LLAMA-2-7B-Chat acts like a gentle reminder. This distinction can be explained by the relatively weak alignment of Mistral-7B-Instruct-v0.2 when compared with LLAMA-2-7B-Chat. We also showcase more DPPs in Appendix H.
Perplexity [$\downarrow$]
Self-Reminder 298.39
Goal Prioritization 40.65
System Prompt 25.65
RPO 8780.94
DPP (Ours) 56.57
Quantitatively, we measure the perplexity for our DPP as well as other defense baseline prompts on LLAMA-2-7B-Chat in Table 6. The perplexity score for a sentence is calculated by averaging the negative log probabilities of next-token, predicted by the GPT-4 model, and using this average as the exponent in a base-2 exponential function. Our method exhibits a lower perplexity score than RPO and Self-Reminder, indicating higher interpretability. It is noteworthy that RPO has the highest perplexity, suggesting that the suffix prompt generated by RPO is highly uninterpretable due to the use of GCG Attack algorithm. Although both Goal Prioritization and System Prompts are hand-crafted defense prompts with lower perplexity (i.e., they are more interpretable prompts), our method remains competitive with these approaches while sparing the need for human interventions in prompt design and optimization.

### v2

@@ 8         Select elite and parent prompts from DPP_Set according to Final Score
10         Find synonyms for each word
11         if random value $<$ WordDict[$synonym$] / sum($word~{}scores$then
12              Replace word with synonym
13         end if
14     end for
15     for $iteration$ in paragraph-level iterations do
16         Repeat line 6 to 8
17         Conduct crossover and mutation on selected parent prompts using Alg. 4
18     end for
@@ 16         Repeat line 6 to 8
21end while
22return (New_DPP_Set, Best_DPP)
@@
Algorithm 1 identifies the optimal DPP for a given pair of refusal and helpful data. Our primary objective is to find a DPP that generalizes well across different user queries. To enhance the universality of DPP, we incorporate $N$ pairs of refusal and helpful data, sampled from their respective datasets. In each iteration of the DPP algorithm, as described earlier, a set of candidate DPPs is generated along with the best DPP for the specific data pair. This set of candidate DPPs is then used for the next pair of refusal and helpful data. By iteratively optimizing this set of DPP candidates, we aim to identify the most generalizable DPP with the best defensive and utility performance. The overall optimization procedure is detailed in Algorithm 5. Appendix C included implementation details and hyperparameter settings.
@@ 4 Experiments
We demonstrate the performance of our DPP through two perspectives: Robustness to standard (non-adaptive) and adaptive jailbreak attacks, Generalization to unforeseen jailbreak queries and different LLMs, and Clarity of the best-found DPPs. All final DPPs are listed in Appendix E.
@@ 4.1 Experimental Setup
Adversarial Dataset: We use the AdvBench (Zou et al., 2023), specifically the harmful behavior instructions  [^2] https://github.com/llm-attacks/llm-attacks/blob/main/data/advbench/harmful_behaviors.csv, as jailbreak questions. Each of them is fed into a well-aligned LM (Llama-2-7B-Chat (Touvron et al., 2023)) to generate the denial responses. In our experiment, we sampled 100 jailbreak questions and recorded them with their refusal responses to create the Adversarial Dataset.
@@
Language Models: We perform our jailbreak experiments on two specific LLMs: Llama-2-7B-Chat (Touvron et al., 2023) and Mistral-7B-Instruct-v0.2 (Jiang et al., 2023). Llama-2-7B-Chat model is an adapted version of Llama-2-7B, specifically configured for chat-based interactions. Mistral-7B-Instruct-v0.2 model is a fine-tuned chat version of Mistral-7B-v0.2. This model demonstrates a stronger ability in performance, outperforming Llama-2-13B model on all benchmarks while maintaining proficiency in language tasks.
@@
Jailbreak Attack Methods: We use several existing jailbreak attack methods to generate advanced malicious prompts. Specifically, for each malicious behavior statement, we apply several different types of jailbreaking attacks: (i) Uninterpretable Jailbreak Attacks – we used GCG (Zou et al., 2023) and Base64 (Wei et al., 2023a) to generate adversarial prompts. Specifically, GCG is used to generate an adversarial suffix for each malicious query. Base64 encodes each harmful query in Base64 format. (ii) Interpretable Jailbreak Attacks – AutoDAN (Liu et al., 2023), PAIR (Chao et al., 2023), TAP (Mehrotra et al., 2023), and ICA (Wei et al., 2023b) are interpretable attacks that we used to translate the original malicious query into a new improved malicious query. Please refer to Appendix F for more details on generating new malicious queries. (iii) Generation-based Jailbreak Attacks – we follow Catastrophic Attack (Huang et al., 2023) to vary the hyperparameters of the LLM to generate malicious responses for each harmful question. In our evaluation, similar to the Adversarial Dataset, we utilize 100 harmful behavior questions from AdvBench to generate new malicious queries [^4] For PAIR and TAP adaptive attacks, we directly utilize the dataset provided in their code-base, which they sample 50 harmful behaviors from AdvBench., all of which will be employed in our experiments.
@@
Jailbreak Defense Methods: We compare our DPP to Self-Reminder (Xie et al., 2023) and Goal Prioritization (Zhang et al., 2023b). They are prompt-based defenses that add defense prompts as a prefix or suffix. For the Llama-2-7B chat model, we also include another defensive suffix approach called RPO (Zhou et al., 2024). For Mistral-7B-Instruct-v0.2, instead of using RPO as a baseline, we compare the results with Plain (Default) System Prompt (Zheng et al., 2024a). We defer the discussion of our choices of baselines for the two LLMs to Appendix G. Additionally, the prompts for each defense baselines can be found in Appendix H.
@@
Evaluation Metrics: We use the Attack Success Rate (ASR) as our primary metric for evaluating the effectiveness of jailbreak defenses. The ASR measures the proportion of malicious queries that successfully bypass the LLMs alignment and generate harmful responses. Details on how we calculate ASR can be found in Appendix I. In addition to ASR, we also use AlpacaEval (Li et al., 2023) to evaluate the utility degradation of the LLM model when defenses are employed. Specifically, we utilize the metric called Win-Rate. This involves comparing the frequency with which outputs from LLM are favored over those from a reference model, given a specific user instruction. Utilizing simulated Win-Rate offers a straightforward, comparable metric across various LLMs using the same reference model. In Appendix J, we discuss the setups of evaluating with Win-Rate.
@@ 4.2 Robustness against Non-adaptive and Adaptive Attacks
RPO (Zhou et al., 2024) 0.000 0.420 0.280 0.190 0.060 0.060 0.168 79.23
Goal Priorization (Zhang et al., 2023b) 0.000 0.020 0.520 0.020 0.020 0.020 0.100 34.29
Self-Reminder (Xie et al., 2023) 0.030 0.290 0.000 0.040 0.020 0.000 0.063 64.84
@@
Table 3: Adaptive Attack Success Rates Rate on Llama-2-7B-Chat model. Our method has the lowest Average ASR.
Self-Reminder 0.410 0.263 0.210 0.080 0.040 0.060 0.177
RPO 0.360 0.653 0.920 0.170 0.240 0.400 0.457
Goal Prioritization 0.660 0.0033 0.190 0.530 0.060 0.040 0.247
DPP (Ours) 0.160 0.247 0.120 0.110 0.080 0.060 0.130
@@
Adaptive Attacks: Adaptive attack (Tramer et al., 2020) is a critical evaluation procedure for assessing defense effectiveness when the defense mechanism is known to the attack. In this study, we assume that the attacker can query the protected large language model (LLM) while defense mechanisms are active during jailbreak attempts. By "adaptive", we refer to the attacker’s ability to target an LLM equipped with a DPP without prior knowledge of the specific DPP being utilized (i.e., DPP is part of the post-query system prompt used by a closed-sourced LLM service provider to improve safety). In this setup, we adapted the attack strategies described in Appendix K. Due to the known limited effectiveness of PAIR and TAP in the non-adaptive setting on the Llama-2-7B-Chat model, (Chao et al., 2023; Mehrotra et al., 2023), we introduce a new adaptive attack: Catastrophic Adaptive Attack. In addition, Base64 attack is a static approach, so the adaptive setting cannot be directly applied to it. Therefore, we remove Base64 attack from the evaluation. Table 3 shows the adaptive attack results. Our method still has the best adaptive ASR with respect to ICA and GCG adaptive attacks. Although Goal Prioritization has the best ASR under catastrophic attacks, which is 0.33%, it fails to defend against ICA and AutoDAN adaptive attacks. On the other hand, our method outperforms Self-Reminder against all adaptive attacks except AutoDAN. Notably, our method attains the best Average ASR, which is 13.0% (outperforming the second-best method by more than 4%), while RPO has the worst robustness, with an Average ASR of 45.7%. In addition to evaluating ASR through keyword-based detection, we also assess it using an Llama-Guard-as-a-judge (Inan et al., 2023) approach. Table 18 in Appendix L illustrates that our DPP outperforms other baseline models, aligning with the findings from the keyword-based evaluation. In Appendix M, we also conducted our DPP with different initialized prototypes and found the defensive performance was consistent. A similar pattern emerges when applying our DPP to defend against two other recent jailbreak attacks, as detailed in Appendix N. In Table 24, DPP achieves 0.0% average ASR in defending against these attacks. In conclusion, non-adaptive and adaptive evaluations affirm that our DPP consistently surpasses other defense mechanisms in robustness, with minimal utility degradation across the board. This comprehensive performance solidifies our method’s position as a preferable choice for defending the model against diverse and sophisticated attacks.
@@ 4.3 Generalization of DPP
Table 4: Attack Success Rates (ASRs) and Win-Rates (utility) on Mistral-7B-Instruct-v0.2 model across six different jailbreak attacks. Our method can achieve the lowest Average attack success rate with reasonable trade-off of Win-Rate when compared with other baselines.
@@
Self-Reminder (Xie et al., 2023) 0.550 0.270 0.510 0.880 0.420 0.260 0.482 88.82
System Prompt (Zheng et al., 2024a) 0.740 0.470 0.300 0.970 0.500 0.180 0.527 84.97
Goal Priorization (Zhang et al., 2023b) 0.030 0.440 0.030 0.390 0.300 0.140 0.222 56.59
@@
Table 5: Adaptive Attack Success Rates on Mistral-7B-Instruct-v0.2. Our method has the lowest Average ASR.
@@
Regarding the trade-off between defense effectiveness and utility degradation, unlike the Llama-2-7B-Chat results, our method exhibits a higher utility degradation, as indicated by the Win-Rate, compared to Self-Reminder, and System Prompt. Nonetheless, the superior defense performance (a gap greater than 46% in average ASR) of our method justifies this increased utility degradation. It is noteworthy that despite the relatively higher utility impact, our method still shows much less degradation compared to the Goal Prioritization approach. Our result suggests that Mistral-7B-Instruct-v0.2 has a worse defense-utility trade-off than Llama-2-7B-Chat. That is, the cost of making Mistral-7B-Instruct-v0.2 robust to jailbreak attacks on utility is more significant than Llama-2-7B-Chat. We present additional experiments in Appendix O, where we compare our results with another defense baseline and observe similar effects.
Adaptive Attacks: Table 5 demonstrates that our method consistently performs best as a defense mechanism against jailbreak attacks on average. Although our approach is slightly less effective in the GCG Adaptive Attack compared to Goal Prioritization, it exhibits superior defensive capabilities in the AutoDAN, PAIR, and TAP adaptive attacks. Similar to the Llama-2-7B-Chat adaptive experiment, we also consider replacing the keyword-based judge with an Llama-Guard-based approach. Table 19 in Appendix L shows that our DPP achieves an average ASR of 5.4%, which is superior to other baselines. Furthermore, we performed additional experiments on two other jailbreak attacks and recorded the results in Appendix N to assess the performance of our DPP.
Unforeseen Jailbreak Queries: We also test the generalization of each defense by sampling another 100 harmful queries from the AdvBench dataset which are independent from the Adversarial Dataset. Then we utilize these harmful queries to test the performance of our DPP against 4 different jailbreak attacks under adaptive settings. In Table 37, the DPP demonstrates superior performance, achieving the lowest Average ASR of 7.5% on Llama-2-7B-Chat model. This indicates that DPP is the most effective defense mechanism against various jailbreak attacks. Specifically, DPP achieves the lowest ASR in TAP and ICA. Similarly, Table 38 shows DPP, on Mistral-7B-Instruct-v0.2, again outperforms other defense baselines, with an Average ASR of 39.4%. DPP illustrates notable performance in AutoDAN and ICA attacks, suggesting enhanced capability in unexpected scenarios compared to other baselines. We also evaluated our DPP under the same conditions using an Llama-Guard-based judge. The results in Table 20 and Table 21 in Appendix L demonstrate consistency with the findings in Table 37 and 38. Furthermore, an evaluation was conducted using the JailbreakBench Chat dataset (JBC) (Chao et al., 2024), which contains harmful queries distinct from those found in the AdvBench dataset. The results are presented in Table 27 located in Appendix P.
@@
4.4 Clarity of DPP
We explore the clarity of our DPP by presenting our DPPs trained on both Llama-2-7B-Chat and Mistral-7B-Instruct-v0.2 models below. Table 30 demonstrates that both DPPs exhibit greater fluency compared to the baseline, RPO. Notably, the optimized DPP for Mistral-7B-Instruct-v0.2 is particularly explicit in issuing refusals when encountering "defective components". In contrast, the DPP for Llama-2-7B-Chat serves as a reminder to "furnish a thorough response". This difference can be attributed to the comparatively weaker alignment of Mistral-7B-Instruct-v0.2 relative to Llama-2-7B-Chat. A more detailed discussion of this distinction is provided in the Appendix Q. Furthermore, additional DPPs are presented in Appendix E.
@@ 4.5 Ablation Study
We report an ablation study to test the effectiveness of each objective functions mentioned in Sec. 3.2. We summarized the result in Table 6. The study was performed under two specific settings: No Defense setting and No Helpful setting. In No Defense setting, where $\alpha=0$ in Eq. 6 (i.e. only optimized on utility score), the GCG Attack score was 16.0%, and the GCG Adaptive Attack score was 19.0%, with a Win Rate of 72.85%. Conversely, in the No Helpful setting, where $\beta=0$ (i.e. only optimized on defense score), the GCG Attack score decreased to 3.0%, and the GCG Adaptive Attack score to 15.0%, while the Win Rate dropped to 65.34%. These findings suggest that disabling either the helpful or defense component significantly reduces the Attack Success Rate (ASR) or the Win Rate. This underscores the importance of both objectives in achieving the most optimal solution. Additional ablation studies show: (1) DPP is more effective as a suffix than a prefix, (2) DPP enhances both defense and utility regardless of initialization, and (3) Replacing HGA with RLPrompt confirms HGA’s superiority in finding effective DPP. (4) Underscoring importance of and sentence-level synonym substitution for both defense and utility in HGA algorithm. Detailed results are in Appendix B.
5 Conclusion
The proposed Defensive Prompt Patch (DPP) framework presents a scalable and practical prompt-based approach to improving LLM safeguards, addressing critical vulnerabilities exposed by jailbreak attacks while preserving utility. The empirical tests conducted – including Llama-2-7B-Chat and Mistral-7B-Instruct-v0.2 models, 7 jailbreak attack strategies, and several state-of-the-art prompt-based defenses – substantiate that DPP effectively reduces the attack success rate to low levels with minimal impact on model performance. Moreover, the adaptability of DPP to function effectively even on less-aligned models underscores its potential as a universal defensive solution in various LLM models. The clarity property inherent in our DPP opens up a new avenue to infusing and accelerating prompt engineering by human users for enhancing LLM safety alignment. Future work will focus on further refining the DPP’s capabilities specifically, enhancing the utility degradation on less-aligned model while preserving the defense capability.
Limitation
Computational Efficiency and Scalability: The DPP training algorithm, which involves a Hierarchical Genetic Algorithm (HGA), is computationally intensive, which we show our computation cost in Appendix C. The scalability of our approach to larger datasets or more extensive model deployments may be limited by the computational resources required for iterative optimization and evaluation. As model sizes and the volume of data grow, the efficiency of DPP training may need further optimization. Nonetheless, once DPP is trained, the deployment of DPP has high inference efficiency by simply attaching the DPP to the user query.
Cost of Training with DPP: The DPP training algorithm requires a LLM to revise the prototype prompt, and currently, we are using GPT-4 as the revising LLM, therefore, the cost of accessing OpenAI platform is considerable high for this training process. In order to minimize the cost of training, one approach is to replace the GPT-4 with some open-sourced LLMs, which will be the future scope of this work.
Limitations of other defense baselines: We noticed that other defense baselines also contain limitations. For Self-Reminder, we notice this training procedure works poorly on Llama-2-7B-Chat model. Since its well-alignment, it will often refuse to improve upon the defense prompt. For RPO, the main limitation is the training time. RPO adopted the GCG attack training procedure, and thus results a high computational cost for finding the defense suffix. We also observe the inefficient of RPO when defending jailbreak attacks which is discussed in Appendix G. Goal Prioritization is strong defense against GCG attack, but it seems less effective when defending AutoDAN, TAP and PAIR attacks. Moreover, it contains a long in-context learning, which cause the inference time when adding Goal Prioritization increases. From both Llama-2-7B-Chat and Mistral-7B-Instruct-v0.2, we observe the utility degradation is large for Goal Prioritization.
Vulnerability to Modification: Our proposed use case for DPP with open-weight models is primarily intended for model providers. These providers aim to deploy services using open-weight models similarly to how closed-source models are utilized. In this context, DPP can be appended after users submit their queries, enhancing the service’s functionality. Conversely, if users run an open-weight model locally, DPP or any system prompts can be easily removed by malicious actors. Thus, the LLMs will still be vulnerable to the Jailbreak Attacks. Under such context, DPP will not be able to protect the actual safety of the open-weighted model.
Acknowledgment and Funding Statement
Chen Xiong and Tsung-Yi Ho, from the JC STEM Lab of Intelligent Design Automation, are funded by the Hong Kong Jockey Club Charities Trust.
Appendix A Motivation for Human-Readable Defense Prompts
We deliberately keep our Defensive Prompt Patch (DPP) human-readable for three reasons:
  1. 1.
    Clarity and stability of the defense mechanism. As noted in Sec. 4.4 and Appendix Q, readable prompts make the mechanism transparent to developers and auditors, who can then verify that the prompt truly enforces the intended safety objective.
  2. 2.
    Generalization across LLMs. Because prompts are appended at inference time, keeping them model-agnostic and human-legible facilitates deployment on less-aligned or open-weight models (Sec. 4.3) without any architecture-specific tuning.
  3. 3.
    Human-centric design philosophy. Prompt-based defenses such as Self-Reminder, Goal Prioritization and DPP deliberately avoid re-training. A readable patch can be inspected, refined, or even crafted from scratch by non-experts, staying true to that principle of simplicity.
In addition, the survey paper Arrieta et al. (2019) stresses that human-readability is critical for the safety and robustness of AI systems. Transparent artefacts allow stakeholders to identify and mitigate vulnerabilities before they can be exploited by malicious actors, thereby aligning system behaviour with human values and ethical norms.
Appendix B Ablation Studies on DPP algorithm
In this section, we want to report the ablation studies that we have done to further understand our DPP methodology. We conduct several different ablation studies:
  • Ablation study on different objective functions as described in Experiment 4.5.
  • Ablation study on different stability and patching format.
  • Ablation study of the impact of prototype initialization on the performance of our DPP Algorithm.
  • Ablation study on different solver by replacing HGA algorithm with RLPrompt Deng et al. (2022) using the objective functions Eq 6.
  • Ablation study on necessity of sentence-level synonym substitution.
B.1 Ablation study on different objective functions
Table 6 contains the ablation results for No Defense setting and No Helpful setting.
Table 6: Ablation study on masking out different objective functions and evaluate the DPP on ASR and Win-Rate.
Coefficient Settings GCG Attack [$\downarrow$] GCG Adaptive Attack [$\downarrow$] Win Rate [$\uparrow$]
No Defense 0.16 0.19 72.85
No Helpful 0.03 0.15 65.34
B.2 Ablation study on different patching format
We also investigate the stability of DPP and its patching format (i.e., as a prefix or as a suffix to an input query). We independently initialized three distinct sets of defense prompts as prefixes and suffixes and applied the DPP algorithm to each set. Table 7 shows the ASR and Win-Rate under both non-adaptive and adaptive GCG attack scenarios for the Llama-2-7B-Chat model.